# Collective emission and selective-radiance in atomic clouds and arrays coupled to a microring resonator

### v1

@@ II.4 Eigenmodes of the excitation
II.5 Photon emission into the cavity and free space
@@
$\begin{split}R&=-\langle\dot{\hat{e}}\rangle=-\frac{d(\vec{\sigma}^{\dagger}\vec{\sigma})}{dt}=\vec{\sigma}^{\dagger}(i\mathbf{M}^{\dagger}-i\mathbf{M})\vec{\sigma}\\ &=2i\vec{\sigma}^{\dagger}\mathbf{G}_{c}\vec{\sigma}-2i\vec{\sigma}^{\dagger}\mathrm{Im}\{\mathbf{G}_{f}\}\vec{\sigma},\end{split}$ (21)
where $\langle\hat{e}\rangle$ is the total excitation probability in the system and we have used $\mathbf{G}_{c}^{\dagger}=-\mathbf{G}_{c}$ and $\mathbf{G}_{f}^{T}=\mathbf{G}_{f}$. The rate of photons emitted into the cavity mode and the free space modes can be separately calculated as
$R_{c}=2i\vec{\sigma}^{\dagger}\mathbf{G}_{c}\vec{\sigma}$ (22)
$R_{f}=-2\vec{\sigma}^{\dagger}\mathrm{Im}\{\mathbf{G}_{f}\}\vec{\sigma}.$ (23)
Experimentally measured transmission rate at the output of the bus waveguide is directly proportional to the photon emission rate $R_{c}$. Photon emission into free space is typically collected by optics spanning over a finite solid angle. If the angular pattern of photon emission is fixed in time, the measured photon rate is proportional to $R_{f}$.
Although the amplitude associated with each eigenstate decays exponentially in time, the emission rates contain the cross-coupling between all the eigenmodes, leading to more complicated time evolution dynamics. Hence, the rates, as well as the total excitation probability, do not decay purely exponentially. The decay rate of the light emitted varies over time. In this paper, the focus will be on the instantaneous decay rate immediately after the drive is turned off in order to study the initial time behaviors of the system. The decay rates of the excitation into the free space and cavity modes have been defined as
$\gamma_{f}=\frac{R_{f}}{\langle\hat{e}\rangle}\quad\mathrm{and}\quad\gamma_{c}=\frac{R_{c}}{\langle\hat{e}\rangle}.$ (24)
The former is approximately independent of the single atom cooperativity $C_{1}$. This is because, at low intensities, the decay rate into free space is solely dependent on the relative phase and the relative amplitude of the excitation between the atoms. While shifting the mean atomic position in the z-axis and changing the atom-cavity interaction strength could change the absolute excitation probability, it does not change the relative amplitude and phase between the atoms.
II.6 Decay rates of photon emission
@@
$\Gamma_{f}=\frac{-\dot{R}_{f}}{R}\quad\mathrm{and}\quad\Gamma_{c}=\frac{-\dot{R}_{c}}{R},$ (25)
@@
$\Gamma_{\mathrm{exp}}=-\frac{\dot{R}_{c}}{R_{c}}=\Gamma_{c}+\Gamma_{f}\theta,$ (26)
where $\theta=(\dot{R}_{c}/R_{c})/(\dot{R}_{f}/R_{f})$.
@@
III Results
We consider two main configurations: atoms in a randomly distributed cloud and an ordered array of atoms. The focus of this study will be on the decay characteristics due to emission into free space and into the cavity.
@@ III Results
The atoms can be excited using either a short Gaussian pulse or a long-duration pulse that drives the system to the steady state. The collective excited states associated with these situations are considerably different. In the case of the short pulse, the atoms will get excited into a state similar to a spin-wave excitation known as the timed-Dicke state (TDS) Scully and Svidzinsky (2009) with a wavenumber corresponding to that of the resonator mode. Since the pulse time is short, the detuning will also play a negligible role. This will be a superradiant state of the atom-cavity interaction. For decay into free space, the TDS has a higher wavenumber than $k_{0}$, and is typically associated with subradiance.
The system can also be driven into the steady state (SS) by using a significantly longer pulse duration. Since there is constant driving, the system is usually pumped into excited states exhibiting slower decaying modes. The detuning also plays an important role in determining which modes are excited in this case.
III.1 Evanescent coupling
The atoms couple with the resonator mode through an evanescent coupling. The coupling decreases exponentially with distance from the resonator. The atoms’ average position along the z-axis depends on the trapping potential, allowing control of the single atom cooperativity $C_{1}$. This provides an opportunity to study the behavior of collective photon emissions as a function of the atom-cavity interaction strength.
The experiment controls the average position of the atoms to be $\geq 400\mathrm{nm}$ away from the surface, which corresponds to $C_{1}\lesssim 0.05$. In the proposed plan to implement an atom array, the trapping methodology can afford the atoms to be much closer to the resonator waveguide, leading to much higher average cooperativities ($C_{1}\sim 10$).
III.2 Atom Cloud
@@
Figure 3: The measured and calculated linewidth of the steady state transmission spectrum as a function of the number of atoms for $C_{1}=0.05$. Blue squares with error bars denote experimental data. Red circles with lines show theoretical calculations. The black triangles and orange diamonds with dashed lines show the theoretical calculation of hypothetical situations where there is no free space collective dipole-dipole interaction or when there are no stochastic effects
@@
In the current state of the experiment, a cloud of atoms is trapped and cooled near the microring resonator in a trap with a cigar-shape of around $2\mu\mathrm{m}$ along the waveguide. In the transverse direction, the cloud is tightly trapped to an r.m.s. of around $100\mathrm{nm}$. The position dependence on the z-direction above the resonator is not Gaussian Zhou et al. (2024). It has an approximate r.m.s. of $430\mathrm{nm}$.
The system can be directly driven with a large wavenumber. Unlike in free space, where systems are conventionally only driven by the free space modes with wavenumber $k_{0}$, the effective refractive index of the resonator waveguide determines the wavenumber of the drive. Figure 2 (a) and (b) explore how driving an atomic cloud with a different wavenumber can affect the free space decay rate. The system is being excited using either short or long pulses, which can exhibit contrasting subradiant (SS) or superradiant (TDS) emission dynamics.
When driven using a short pulse, the system is excited into the TDS with the wavenumber $k\sim 1.7k_{0}$ While one might expect this state to be subradiant to free space, Ref. Grava et al. (2022) has shown that for an atom cloud, the smallest decay rate for such high-k states is $\Gamma_{0}$. This can be seen in Fig. 2 (b), where the free space decay rate $\gamma_{f}$ for all atom numbers converge to $\Gamma_{0}$ as the wavenumber increases beyond $k_{0}$. This happens because the dephasing from atoms being very close to each other in a random cloud acts against the longer-range coherence that is built up to cause subradiance. The TDS will be perfectly superradiant to the cavity with an average decay rate of $NC_{1}\Gamma_{0}$ while having a decay rate of $\Gamma_{0}$ into free space. Therefore, the TDS will have a total decay rate of $(1+NC_{1})\Gamma_{0}$.
@@
These effects can also be seen in Fig. 2 (c) and (d), which show the distribution of the decay rates into the free space and cavity modes without ensemble averaging. The free space decay rate ($\gamma_{f}$) is suppressed in the case of the SS while clustered around $\Gamma_{0}$ for the TDS. For the decay rate into the cavity ($\gamma_{c}$), the TDS has a large spread of decay rates, but the average is close to $NC_{1}\Gamma_{0}$, where $N$ is the atom number in the system. On the other hand, for the SS, $\gamma_{c}$ averages to around $20C_{1}\Gamma_{0}$, which is still superradiant but not as good as the with the TDS.
The large spread in $\gamma_{c}$ is primarily due to the stochastic non-uniformity of the single atom cooperativity $C$, as mentioned in Sec. III.1. When $C$ is uniform for all the atoms, the distribution of $\gamma_{c}$ will be a delta function at $NC_{1}\Gamma_{0}$ for the TDS and will have a considerably smaller spread for the SS. This emphasizes the importance of minimizing the spread in the z-position for systems that have evanescent coupling.
On the other hand, the steady state transmission spectrum through the output bus waveguide can be calculated and compared with experimental measurements. The linewidth can provide some indication of the the decay properties of the system. Figure 3 compares the experimentally measured and calculated linewidths of the spectrum, which are both determined using simple Lorentzian fits. The mean atom number $N$ in the experiment is extracted by fitting the steady-state transmission spectrum assuming a Gamma distribution of $NC$ with mean value $C_{N}$ and $N=C_{N}/C_{1}$ (see Ref. Zhou et al. (2024)).
@@
However, in practical implementations, the expected linewidth suppression is significantly weakened due to multiple stochastic effects. The primary contributor is the non-uniformity in single-atom cooperativity $C$ (see Sec. III.1), which broadens the linewidth. Additionally, the AC Stark shift from trapping fields varies with atomic position, introducing a slight spectral asymmetry and further broadening the linewidth. Moreover, the number of trapped atoms follows a Poissonian distribution, which introduces further broadening, especially for small $N$. The orange diamonds in Fig. 3 depict the hypothetical where these stochastic effects are absent. This linewidth shows a greater degree of suppression for larger $N$ and approaches $\Gamma_{0}$ when $N\rightarrow 1$.
@@
In the associated experimental paper Zhou et al. (2025), the ratio $\theta$ in Eq. (26) serves as an important factor to correlate experimentally measured decay rates to the actual decay rates of the system. To get an intuition behind this $\theta$, we can express it as $\theta=(\Gamma_{c}/\Gamma_{f})/(\gamma_{c}/\gamma_{f})$. In the case of the TDS, $\Gamma_{c}/\Gamma_{f}<NC_{1}$ resulting in the $\theta<1$, while for the SS, $\Gamma_{c}/\Gamma_{f}>NC_{1}$ resulting in the $\theta>1$. Further analysis on the implications of $\theta$ is explored in the Supplementary of Ref. Zhou et al. (2025).
III.3 Atom Arrays
@@
Although a disordered atomic ensemble can become subradiant to free space through a steady state excitation, the degree of superradiant interaction with the cavity will also decrease, resulting in $\gamma_{c}/\gamma_{f}\sim NC_{1}$, similar to the superradiant scaling found in the TDS. The TDS, on the other hand, is more useful in utilizing the full superradiance benefit of this platform. In this section, we explore the potential of atom arrays in reaching subradiance in free space through the simple TDS excitation.
The next stage of the experiment involves the implementation of an array of atoms trapped above the ring resonator. According to Chang et al. (2019), the atoms will be trapped using two-color evanescent fields, resulting in the atoms being separated by about $0.3\lambda$, where $\lambda$ is the resonant wavelength of the light. This is close to half the wavelength of the high-k excitation mode of the ring resonator. When the atoms are excited through the resonator, it can facilitate almost perfectly subradiant states into free space due to the alternating phases in adjacent atoms.
The higher wavenumber of the waveguide excitation $k_{\mathrm{wg}}$ plays a major role in contributing to accessing highly subradiant states in the case of atom arrays. Experimentally, this is only dependent on the effective refractive index $n_{\mathrm{eff}}=k_{\mathrm{wg}}/k_{0}$ of the WGM. The decay rate into free space as a function of $k_{\mathrm{wg}}$ and the separation of the atom $d$ has been depicted in Fig. 5 for a perfect atom array with no disorder. Although very good subradiance ($\gamma_{f}<0.1\Gamma_{0}$) can only be achieved with $n_{\mathrm{eff}}>1$ and separations $d<0.5\lambda$, there are still other regions where nominal subradiance ($\gamma_{f}<\Gamma_{0}$) can be achieved. As mentioned at the end of Sec. II.5, in the low-intensity limit, $\gamma_{f}$ is only dependent on the relative phase between the atomic excitation and is therefore independent of the single atom cooperativity $C_{1}$.
The white dot in Fig. 5 marks the proposed parameters of the upcoming experiment. For a perfect array with $N=20$, separation of $d=0.3\lambda$, and effective refractive index $n_{\mathrm{eff}}=1.69$, the free space decay rate $\gamma_{f}=0.035\Gamma_{0}$.
III.3.1 Effects of Disorder
@@ III.3.2 Circular Array
Figure 7 shows the comparison between a linear array and a ring-shaped array trapped along the microring resonator. It depicts the decay rate into free space that is achievable using the TDS excitation through the waveguide as a function of the number of atoms in the array. This is similar to Fig. 7 in Ref. Asenjo-Garcia et al. (2017). We can almost reach a perfect subradiant state with the ring resonator setup and evanescent field traps.
@@
Although there is great promise with a ring-shaped array due to the exponential scaling, the effects of disorder described in Sec. III.3.1 will still limit the effectiveness of the system.
IV Conclusion

### v2

@@ III.1 Steady state vs timed-Dicke state
III.2 Photon emission into the cavity and free space
@@
$\begin{split}R(t)&=-\langle\dot{\hat{e}}\rangle=-\frac{d(\vec{\sigma}^{\dagger}\vec{\sigma})}{dt}=\vec{\sigma}^{\dagger}(i\mathbf{M}^{\dagger}-i\mathbf{M})\vec{\sigma}\\ &=2i\vec{\sigma}^{\dagger}\mathbf{G}_{c}\vec{\sigma}-2i\vec{\sigma}^{\dagger}\mathrm{Im}\{\mathbf{G}_{f}\}\vec{\sigma},\end{split}$ (23)
where $\langle\hat{e}\rangle$ is the total excitation probability in the system, and we have used $\mathbf{G}_{c}^{\dagger}=-\mathbf{G}_{c}$ and $\mathbf{G}_{f}^{T}=\mathbf{G}_{f}$. The time dependence of $\langle\hat{e}\rangle$ and $\vec{\sigma}$ is not explicitly written for the sake of simplicity. The rate of photons emitted into the cavity mode and the free space modes can be separately calculated as
$R_{c}(t)=2i\vec{\sigma}^{\dagger}\mathbf{G}_{c}\vec{\sigma}$ (24)
$R_{f}(t)=-2\vec{\sigma}^{\dagger}\mathrm{Im}\{\mathbf{G}_{f}\}\vec{\sigma}.$ (25)
The experimentally measured transmission rate at the output of the bus waveguide is directly proportional to the photon emission rate $R_{c}$. Photon emission into free space is typically collected by optics spanning over a finite solid angle. If the angular pattern of photon emission is fixed in time, this measured photon rate is proportional to $R_{f}$.
Although the amplitude associated with each eigenstate decays exponentially in time, the emission rates contain the interference between all the eigenmodes, leading to more complicated time evolution dynamics. Hence, the total excitation probability does not decay purely exponentially. Correspondingly, the photon emission rates, as well as the decay rates, become time-dependent. In this paper, the focus will be on the instantaneous decay rate immediately after the drive is turned off in order to study the initial time behaviors ($t\ll 1/\Gamma_{0}$) of the system. The decay rates of the excitation into the free space and cavity modes have been defined as
$\gamma_{f}=\frac{R_{f}(0)}{\langle\hat{e}\rangle(0)}\quad\mathrm{and}\quad\gamma_{c}=\frac{R_{c}(0)}{\langle\hat{e}\rangle(0)}.$ (26)
The former is approximately independent of the atom-cavity interaction strength, i.e., the average single atom cooperativity $C_{1}$. This is because, at low intensities, the decay rate into free space is solely dependent on the relative phase and the relative amplitude of the excitation between the atoms. While shifting the mean atomic position in the z-axis and changing the atom-cavity interaction strength could change the absolute excitation probability, it does not change the relative amplitude and phase between the atoms.
III.2.1 Decay rates of photon emission
@@
$\Gamma_{f}=\frac{-\dot{R}_{f}(0)}{R(0)}\quad\mathrm{and}\quad\Gamma_{c}=\frac{-\dot{R}_{c}(0)}{R(0)},$ (27)
@@
$\Gamma_{\mathrm{exp}}=-\frac{\dot{R}_{c}(0)}{R_{c}(0)}=\Gamma_{c}+\Gamma_{f}\theta,$ (28)
where $\theta=[\dot{R}_{c}(0)/R_{c}(0)]/[\dot{R}_{f}(0)/R_{f}(0)]$.
@@
IV Results
We consider two main configurations: atoms in a randomly distributed cloud and an ordered array of atoms. We explore the potential of both these configurations in utilizing collective states to suppress emission into free space while promoting stronger coupling to the cavity mode. The primary focus will be on comparing the decay characteristics due to emission into free space and the cavity.
@@ IV Results
IV.1 Atom Cloud
@@
Figure 3: The measured and calculated linewidth of the steady state transmission spectrum as a function of the number of atoms for $C_{1}=0.05$. Blue squares with error bars denote experimental data. Red circles with lines show theoretical calculations. The black triangles and orange diamonds with dashed lines show the theoretical calculation of hypothetical situations where there is no free space collective dipole-dipole interaction or when there are no stochastic effects.
@@
In the current state of the experiment, a cloud of atoms is trapped and cooled near the microring resonator in a trap with a cigar-shape with an r.m.s. of around $2\mu\mathrm{m}$ along the waveguide. In the transverse direction, the cloud is tightly trapped to an r.m.s. of around $100\mathrm{nm}$. In the z-direction, the cloud has an approximate r.m.s. of $430\mathrm{nm}$, but has an asymmetric density profile due to the shape of the trap, as detailed in Ref. Zhou et al. (2024).
The system can be resonantly driven ($\omega=\omega_{A}$) with a large wavenumber ($k>k_{0}$). Unlike in free space, where systems are conventionally only driven by the free space modes with wavenumber $k_{0}$, the effective refractive index of the resonator waveguide determines the wavenumber of the drive. By increasing the excitation wavenumber, the rapid spatial variation in phase allows the system to access different collective excited states with varied decay rates.
Figure 2 (a) and (b) explore how driving an atomic cloud with a different wavenumber can affect the free space decay rate. The system is being excited using either short or long pulses, which can exhibit contrasting subradiant (SS) or superradiant (TDS) emission dynamics.
When driven using a short pulse, the system is excited into the TDS with the wavenumber $k\sim 1.7k_{0}$, where $n_{\mathrm{eff}}\approx 1.7$ is numerically obtained based on the waveguide geometry presented in Ref. Zhou et al. (2024, 2025). Although this state will be subradiant to free space in ordered arrays, Ref. Grava et al. (2022) has shown that for an atom cloud, the smallest decay rate for such high-k states is $\Gamma_{0}$. This can be seen in Fig. 2 (b), where the free space decay rate $\gamma_{f}$ for all atom numbers converge to $\Gamma_{0}$ as the wavenumber increases beyond $k_{0}$. This happens because the dephasing from atoms being very close to each other in a random cloud acts against the longer-range coherence that is built up to cause subradiance. The TDS will be perfectly superradiant to the cavity with an average decay rate of $NC_{1}\Gamma_{0}$ while having a decay rate of $\Gamma_{0}$ into free space. Therefore, the TDS will have a total decay rate of $(1+NC_{1})\Gamma_{0}$.
@@
These effects can also be seen in Fig. 2 (c) and (d), which show the distribution of the decay rates into the free space and cavity modes without ensemble averaging. The free space decay rate ($\gamma_{f}$) is suppressed in the case of the SS, while clustered around $\Gamma_{0}$ for the TDS. For the decay rate into the cavity ($\gamma_{c}$), the TDS has a large spread of decay rates, but the average is close to $NC_{1}\Gamma_{0}$, where $N$ is the atom number in the system. On the other hand, for the SS, $\gamma_{c}$ averages to around $20C_{1}\Gamma_{0}$, which is still superradiant but not as good as in the case with the TDS. This shows that the TDS is better for utilizing superradiant effects.
The large spread in $\gamma_{c}$ is primarily due to the stochastic non-uniformity of the single atom cooperativity $C$, as mentioned in Sec. II.2.1. Instead, in the hypothetical situation where the atoms have no z-spread and the $C$ is uniform for all the atoms, the distribution of $\gamma_{c}$ will be a delta function at $NC_{1}\Gamma_{0}$ for the TDS, and will have a considerably smaller spread for the SS. This direct connection between the broadening of the spectrum and the spread in z-position emphasizes the importance of tight confinement for systems that have evanescent coupling.
On the other hand, the steady state transmission spectrum through the output bus waveguide can be calculated and compared with experimental measurements. The linewidth can provide some indication of the decay properties of the system. Figure 3 compares the experimentally measured and calculated linewidths of the spectrum, which are both determined using simple Lorentzian fits. The mean atom number $N$ in the experiment is extracted by fitting the steady-state transmission spectrum assuming a Gamma distribution of $NC$ with mean value $C_{N}$ and $N=C_{N}/C_{1}$ (see Ref. Zhou et al. (2024)).
@@
However, in practical implementations, the expected linewidth suppression is significantly weakened due to multiple stochastic effects. The primary contributor is the non-uniformity in single-atom cooperativity $C$ (see Sec. II.2.1), which broadens the linewidth. Additionally, the AC Stark shift from trapping fields varies with atomic position, introducing a slight spectral asymmetry and further broadening the linewidth. Moreover, the number of trapped atoms follows a Poissonian distribution, which introduces further broadening, especially for small $N$. The orange diamonds in Fig. 3 depict the hypothetical, where these stochastic effects are absent. This linewidth shows a greater degree of suppression for larger $N$ and approaches $\Gamma_{0}$ when $N\rightarrow 1$.
@@
In the associated experimental paper Zhou et al. (2025), the ratio $\theta$ in Eq. (28) serves as an important factor to correlate experimentally measured decay rates to the actual decay rates of the system. To get an intuition behind this $\theta$, we can express it as $\theta=(\Gamma_{c}/\Gamma_{f})/(\gamma_{c}/\gamma_{f})$. In the case of the TDS, $\Gamma_{c}/\Gamma_{f}<NC_{1}$ resulting in the $\theta<1$, while for the SS, $\Gamma_{c}/\Gamma_{f}>NC_{1}$ resulting in the $\theta>1$. Further analysis on the implications of $\theta$ is explored in the Supplementary Material of Ref. Zhou et al. (2025).
IV.2 Atom Arrays
@@
Although a disordered atomic ensemble can exhibit subradiance to free space under steady-state excitation, the effect is limited. In contrast, ordered arrays can mitigate randomness and promote the coherence necessary for pronounced subradiance. In this section, we explore the potential of such atom arrays in reaching subradiance in free space through the simple TDS excitation.
The next stage of the experiment involves the implementation of an array of atoms trapped above the ring resonator as depicted in Fig. 1 (c). According to Chang et al. (2019), the atoms can be trapped using two-color evanescent fields, resulting in the atoms being separated by about $0.3\lambda_{0}$, where $\lambda_{0}$ is the resonant wavelength of the light in free space. This separation is approximately half the wavelength of the high-wavenumber guided mode in the ring resonator.
When the atoms are excited through the resonator, it can facilitate almost perfectly subradiant states into free space due to the alternating phases in adjacent atoms.
The higher wavenumber of the waveguide excitation $k_{\mathrm{wg}}$ plays a major role in contributing to accessing highly subradiant states in the case of atom arrays. Experimentally, this is only dependent on the effective refractive index $n_{\mathrm{eff}}=k_{\mathrm{wg}}/k_{0}$ of the WGM. The decay rate into free space as a function of $k_{\mathrm{wg}}$ and the separation of the atom $d$ has been depicted in Fig. 5 for a perfect atom array with no disorder. Although very good subradiance ($\gamma_{f}<0.1\Gamma_{0}$) can only be achieved with $n_{\mathrm{eff}}>1$ and separations $d<0.5\lambda_{0}$, there are still other regions where nominal subradiance ($\gamma_{f}<\Gamma_{0}$) can be achieved. As mentioned at the end of Sec. III.2, in the low-intensity limit, $\gamma_{f}$ is only dependent on the relative phase between the atomic excitation and is therefore independent of the average single atom cooperativity $C_{1}$.
The white dot in Fig. 5 marks the proposed parameters of the upcoming experiment. For a perfect array with $N=20$, separation of $d=0.3\lambda_{0}$, and effective refractive index $n_{\mathrm{eff}}=1.69$, the free space decay rate $\gamma_{f}=0.035\Gamma_{0}$. The atoms are assumed to be trapped around $330\mathrm{nm}$ from the resonator to give an average $C_{1}=0.05$. If the atoms are trapped much closer, as is proposed to be possible by Ref. Chang et al. (2019), much larger cooperativities can be achieved. Modeling this would require modification to the free space Green’s function due to the vicinity of the dielectric, as well as going beyond the adiabatic elimination approximation.
IV.2.1 Effects of Disorder
@@ IV.2.2 Circular Array
Figure 7 shows the comparison between a linear array and a ring-shaped array trapped along the microring resonator. It depicts the decay rate into free space that is achievable using the TDS excitation through the waveguide as a function of the number of atoms in the array. The comparison considers atoms trapped in arrays with $0.3\lambda_{0}$ spacing, arranged in either a line or a ring above a hypothetical microring resonator that follows the array geometry, with uniform coupling at $C_{1}=0.05$. This comparison allows us to explore how the system scales with increasing atom number. In a practical setting, such a microring resonator could support a ring of about 400 atoms at the given spacing. This is similar to Fig. 7 in Ref. Asenjo-Garcia et al. (2017). We can almost reach a perfect subradiant state with the ring resonator setup and evanescent field traps.
@@
Although there is great promise with a ring-shaped array due to the exponential scaling, the effects of disorder described in Sec. IV.2.1 will still limit the effectiveness of the system.
V Conclusion